[figure: 9.7 cm]

INFLUENCE OF Al THICKNESS ON MAGNETISATION PROCESS OF Co/Al-O/(Ni_{80}Fe_{20}) TRILAYERS

T. Luciński^{a,b}, S. Czerkas^{b}, H. Brueckl^{b}, G. Reiss^{b}

aInstitute of Molecular Physics, Polish Academy of Sciences, PL 60-179 Poznań,

Smoluchowskiego 17, Poland

bFaculty of Physics, University of Bielefeld, D-33501 Bielefeld, Universitaet Str. 25, Germany

We have investigated the magnetisation reversal processes of Co/Al-O/(Ni_{80}Fe_{20}) heterostructures as a function of Al thickness (d_{Al}). The Co(25nm)/Al(d_{Al})-O/(Ni_{80}Fe_{20}) (25nm) trilayers with 0.5 nm≤ d_{Al}≤ 5 nm were deposited onto SiO_{x} wafers by evaporation in UHV system. The Al spacer layers were naturaly oxidised in air before deposition of the top (Ni_{80}Fe_{20}) layer. The magnetic hysteresis loops were measured in-situ during deposition process as a function of thickness as well as a function of time during oxidation process. Our magnetisation measurements revealed the existence of the critical Al thickness d_{cAl} ≈ 1.2nm. For d_{Al} < d_{cAl} a direct ferromagnetic coupling between Co and Ni_{80}Fe_{20} was observed, however for d_{Al}>d_{cAl} well separated magnetisation loops were found. The in-situ conductance measurements allowed us to evaluate a thickness of aluminium oxide d_{Al−O} formed during oxidation process. It has been found that after 24 hours of oxidation d_{Al−O}≅ 0.5 nm independently on d_{Al}. For selected samples the ex-situ atomic force microscopy measurements have been performed after each step of deposition process in order to check the surface (interface) roughness of the already deposited materials.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : poster

Corresponding author : T. Lucinski

Address for correspondence : Institute of Molecular Physics Polish Academy of Sciences PL-60-179 Poznan, Smoluchowskiego 17, Poland

Email address : lucinski@ifmpan.poznan.pl